A mass is oscillating with amplitude A at the end of the spring. How far is this mass from the equilibrium if the elastic potential energy equals γ times the kinetic energy?


1
Expert's answer
2021-06-03T13:54:26-0400

Mechanical energy is conserved

K+U=E

"E=\frac{1}{2}kA^2" and "U=\frac{1}{2}kx^2"

E=U

"\frac{1}{2}kA^2=\frac{1}{2}kA^2 \implies x=\frac{x}{\sqrt{2}}"